My mothers age is three times my age. The sum of our ages is 40. How old am I? How old is my mother?
Answer by ilana(307)

Say your age is x and your mother's age is y. Then you can write 2 equations.
y=3x and x+y=40
So substitute 3x in for y in the second equation.
x+3x=40
4x=40
x=10
Now solve for y.
y=3x=3(10)=30
Your mother is 30.